After graduation, you have been offered an engineering job with a large company that has offices in
Manila and Cebu. The salary is Php 55,000 per year at either location. Manila’s tax burden (local taxes)
is 6% and Cebu’s is 3.07%. If you accept the position in Cebu and stay with the company for 10 years,
what is the FW of the tax savings? Your personal MARR is 10% per year.
1
Expert's answer
2021-10-10T09:33:41-0400

Salary offered = $55,000 per year

Tax rate in Tennessee = 6%

Tax rate in Pennsylvania = 3.07%

So, Tax per year in Tennessee

"= \frac{55,000*6}{100}\\

 = \$3,300"

Tax per year in Pennsylvania

"= \frac{55,000*3.07}{100}\\

 = \$1,688.5"

So, if you accept the position in Pennsylvania then tax savings per year would be

= Tax per year in Tennessee - Tax per year in Pennsylvania

"= \$3,300 - \$1,688.5\\

= \$1,611.5"

So, the tax savings per year would be $1,611.5 for 10 years.

Now using Future worth (FW)/Annuity (A)relation-

"FW = \frac{A}{r}[(1+r)^n-1]\\
FW = \frac{1611.5}{0.1}[(1+0.1)^10-1]\\
FW = 16115*1.5937\\
FW=\$ 25682"